OpenStudy (anonymous):

Last year, the personal best high jumps of track athletes in a nearby state were normally distributed with a mean of 221 cm & a standard deviationof 11 cm. What is the probability that a randomly selected high jumper has a personal best between 199 & 210 cm?

OpenStudy (anonymous):

why don't you try it out baby ;)

OpenStudy (secret66):

1) P(199 < x < 210) = ? z = (199 - 221)/11 = -2 z = (210 - 221)/11 = -1 P(199 < x < 210) = P(-2 < z < -1) = 0.1359 (Ans.)